1. if cos A = -1/5 and A is between 90 and 180 (degrees) , find sin A/2

2. A plane is flying with an airspeed of 180 miles per hour with a heading of 118 (degrees). The wind currents are a constant 28 miles per hour in the direction due north. Find the true course and ground speed of the plane
Answer #1

Using basic principles of trigonometry you can find the solution of question 1 and using the law of parallelogram of vector addition you can easily find the ground speed and true course of plane. Please find the hand written solution for these questions. I hope it will help you to understand. Please like and comment if you found helpful. Thanks and best wishes.

Answer I. 90<A < 180. Los A = - 12 Cos 20= 1-2 sine Cos2(A) = 1- 2 sin? A 2 sin? A = 1- COSA * - COSA » Sin A = -2 Put the vaAnswer 2. speed of 180 mph at 118 180 a 108=8 = 180 on diagram, o represents airspeed of 180 mph AB represents wind current sfrom north True and speed vector course = Angle of ground speed vector By using Law of parallelogram, tano = oA sino OB + OA